# Finitary conditions in special semidirect products

### v1

@@ Abstract.
We initiate the study of finitary conditions for monoids arising as expansions via the semidirect product construction of a monoid $M$ acting on the left of its power set (which we term ‘special semidirect products’, and denote by $\mathcal{S}(M)$). Specifically, we examine the condition of weak left coherence (every finitely generated left ideal has a finite presentation as a left act); the related conditions of property (L), left ideal Howson, finitely left equated; and the left-right dual properties of weak right coherence, property (R), right ideal Howson and finitely right equated. Each of these conditions is preserved under retract, from which it is immediate that if $\mathcal{S}(M)$ satisfies one of our finitary conditions, then so must $M$, but the converse is not true. Indeed, we show that $\mathcal{S}(M)$ satisfies property (L) (or property (R)) if and only if $M$ is finite. We provide a characterisation of the monoids $M$ such that $\mathcal{S}(M)$ is left (respectively, right) ideal Howson, and give sufficient conditions for $\mathcal{S}(M)$ to be finitely left (respectively, right) equated and hence weakly left (respectively, right) coherent.
@@
Expansions, special semidirect products, finitary conditions
@@ 1. Introduction
This paper initiates the study of finitary conditions (that is, conditions satisfied by every finite monoid) in the context of certain expansions of monoids. Monoid expansions provide systematic ways to build a new monoid from a given input monoid, often by inflating the set of idempotents in some structured manner. They were introduced in [1] to facilitate the application of techniques in finite semigroup theory to infinite monoids and semigroups. Most well-studied examples of expansions have the property that when restricting the input to an interesting class of monoids (e.g. groups, cancellative monoids, unipotent monoids), the output of the expansion is a monoid with desirable structural properties. For example, the Szendrei expansion [18] maps: groups to inverse monoids [1], right cancellative monoids to left ample monoids [5] and unipotent monoids to (what are now called) left restriction monoids [6]. In fact, [1] used an earlier formulation, known as the prefix expansion, which is equivalent to the Szendrei expansion in the case the monoid is a group; [14] investigates the prefix expansion of an arbitrary monoid. The underlying construction used in these expansions is a semidirect product formed from the monoid $M$ acting on a particular semilattice of subsets of $M$. It is then natural to consider the ‘parent’ expansion where the semilattice contains all subsets of $M$, and this is our focus here. Specifically, this paper considers the expansion formed via the action of a monoid $M$ on its power set $\mathcal{P}(M)$ viewed as a monoid (indeed, a semilattice) with respect to the operation of union, and with identity element $\emptyset$, with left action of $M$ on $\mathcal{P}(M)$ given by $(m,X)\mapsto mX=\{mx:x\in X\}$ for all $X\in\mathcal{P}(M)$ and all $m\in M$. Throughout the paper we will write $\mathcal{S}(M)$ to denote $\mathcal{P}(M)\rtimes M$ with product given by $(X,x)(Y,y)=(X\cup xY,xy)$, and with identity element $(\emptyset,1)$, where $1$ is the identity element of $M$, and we shall refer to monoids of this form as ‘special semidirect products’.
@@
Our interest in special semidirect products comes from several directions. First, $\mathcal{S}(M)$ contains the Szendrei expansion of $M$ as a monoid subsemigroup, and (as with the Szendrei expansion) if $G$ is a group then $\mathcal{S}(G)$ is an inverse monoid. If $M=F_{X}$ is the free group on a set $X$, we remark that the monoid $\mathcal{S}(F_{X})$ contains: the free inverse, free ample and free left ample monoids on the same set $X$. For a monoid $M$ which is not a group, the semidirect product $\mathcal{S}(M)$ is no longer inverse (indeed, it is not even regular; see Lemma 3.3 below for details). It is then natural to ask under which circumstances $\mathcal{S}(M)$ belongs to one of several more general classes of monoids (e.g. left/right restriction, ample, adequate, abundant).
@@
Second, we note that any inverse monoid satisfies several finitary conditions, including weak left (respectively, right) coherence [2]. A monoid $M$ is said to be weakly left (respectively, right) coherent if every finitely generated left (respectively, right) ideal of $M$ is finitely presented as a left (respectively, right) $M$-act. A classical result of Wheeler [19], interpreted for acts in [9], says that a monoid is left coherent if and only if the class of existentially closed left acts is axiomatisable. Inspired by this, weak left coherency for monoids, and the related finitary conditions we consider here, have also been formulated in terms of the model-theoretic notion of axiomatisability: see [8, 17, 4]. In previous work [13] the construction $\mathcal{S}(M)$ has been utilised to exhibit some interesting examples of monoids that are neither left nor right coherent. We have remarked that the monoid $\mathcal{S}(F_{X})$ contains the free inverse, free ample and free left ample monoids on set $X$, and it is known that each of the latter is weakly left and right coherent [2, 3, 11]; cf. the table in [13]. One is then drawn to ask (see Question 8.6 of [13]): for which monoids $M$ is $\mathcal{S}(M)$ weakly left (respectively right) coherent? It is known that $M$ is weakly left coherent if and only if it is simultaneously left ideal Howson (the intersection of any two finitely generated left ideals is finitely generated) and finitely left equated (the left annihilator congruences of $M$ are finitely generated as left congruences). We therefore examine the behaviour of these two conditions and the closely related conditions of (R) and (L) with respect to the expansion $\mathcal{S}$.
@@
The paper is structured as follows. In Section 2 we outline the preliminary definitions and results required for our paper and in Section 3 we prove some foundational results on special semidirect products. Our main results are in Section 4. We give precise characterisations of the monoids $M$ for which (R), (L) (Subsection 4.1) or the left and right ideal Howson properties hold in $\mathcal{S}(M)$ (Subsection 4.2). We exhibit some necessary conditions for $\mathcal{S}(M)$ to be finitely right or left equated and provide some configurations of elements in $M$ that prevent $\mathcal{S}(M)$ from being finitely right or left equated (Subsection 4.3). In particular, we apply our results to demonstrate that if $M$ is a right (left) LCM monoid then $\mathcal{S}(M)$ is weakly right (left) coherent. We provide a number of examples and counterexamples to illustrate and separate the properties under consideration. We finish with some open questions.
@@ 2. Preliminaries
In this section we give the necessary specialist background to follow this paper. We refer the reader to [15] for a fuller account of the semigroup theoretic notions we discuss. Throughout the paper $M$ denotes a monoid with identity element $1$.
@@ 2.3. A suite of finitary conditions
Another related finitary condition arises as follows:
@@
It is easy to see that $\mathbf{R}(a,b)$ is a right subact of $M\times M$. We say that $M$ satisfies condition (R) if for all $a,b\in M$ the right subact $\mathbf{R}(a,b)$ is finitely generated.
@@
3. Special semidirect products
@@
    $\mathcal{S}(M)$ is always a proper left $\mathcal{E}$-restriction monoid and so certainly left $\mathcal{E}$-Ehresmann; it is left $\mathcal{E}$-ample if and only if $M$ is right cancellative.
@@
We note that Lemma 3.3 together with Proposition 3.4 yields that the expansion $\mathcal{S}$ maps groups to proper inverse monoids, unipotent monoids to proper left restriction monoids, and unipotent right cancellative monoids to left ample monoids. Similarly, by Proposition 3.5 we have that $\mathcal{S}$ maps right abundant monoids to right abundant monoids and left cancellative monoids to right adequate monoids.
@@ 4.2. The ideal Howson properties
The semidirect product $\mathcal{S}(M)$ is right ideal Howson if and only if $M$ is strongly right ideal Howson. Moreover, $\mathcal{S}(M)$ is principally right ideal Howson if and only if $M$ is principally right ideal Howson.
@@
The semidirect product $\mathcal{S}(M)$ is left ideal Howson if and only if $M$ is left co-ordinated. Further, $\mathcal{S}(M)$ is principally left ideal Howson if and only if $M$ is left $1$-co-ordinated.
@@
Let $M$ be right cancellative. The semidirect product $\mathcal{S}(M)$ is left ideal Howson if and only if $M$ is left ideal Howson.
@@
If $M$ is left (1-)co-ordinated, and $T$ is a retract of $M$, then $T$ is also left (1-)co-ordinated.
@@ 4.3. The finitely equated conditions
Since $M$ is a retract of $\mathcal{S}(M)$ it follows from [4, Corollary 4.12] that a necessary condition for $\mathcal{S}(M)$ to be finitely right (respectively, left) equated is that $M$ is finitely right (respectively, left) equated, however, as we shall see below, the converse does not hold (in either case). We look first at the case of finitely right equated monoids.
Remark 4.15.

### v2

# Subset expansions of monoids

@@ Abstract.
We initiate the study of the expansion $\mathcal{S}(M)$ of a monoid $M$ obtained via the semidirect product of $M$ acting naturally on the left of its power set (regarded as a semilattice under union). We term this the ‘subset expansion’ of $M$. The monoid $\mathcal{S}(M)$ contains the images of several expansions of $M$ of wide interest and use in semigroup theory, in particular the prefix and Szendrei expansions (in the case where $M$ is free, these ‘smaller’ expansions produce free algebras in certain varieties).
We first focus on algebraic properties, specifically those determined by idempotents. We note that $\mathcal{S}(M)$ is regular if and only if $M$ is a group and determine the conditions on $M$ that ensure $\mathcal{S}(M)$ possesses certain generalised regularity properties. Particularly, we show that the expansion $\mathcal{S}$ maps groups to proper inverse monoids, unipotent monoids to proper left restriction monoids, right cancellative monoids to left ample monoids, right abundant monoids to right abundant monoids, and left cancellative monoids to right adequate monoids.
Subsequently, we focus on finitary conditions. We examine the condition of weak left coherence (every finitely generated left ideal has a finite presentation as a left act); the related conditions of property (L), left ideal Howson, finitely left equated, and each of the corresponding left-right dual notions. Each of these conditions is preserved under retract, from which it is immediate that if $\mathcal{S}(M)$ satisfies one of our finitary conditions, then so must $M$, but the converse is not true. For each of our properties to ‘lift’ from $M$ to $\mathcal{S}(M)$ it must undergo a ‘strengthening’. Indeed, we show that $\mathcal{S}(M)$ satisfies property (L) (or its left-right dual) if and only if $M$ is finite. We provide exact characterisations of the monoids $M$ such that $\mathcal{S}(M)$ is: left (or right) ideal Howson; finitely left equated; and (consequently) weakly left coherent. We give sufficient conditions for $\mathcal{S}(M)$ to be finitely right equated and hence weakly right coherent.
@@
Expansions, semidirect products, finitary conditions
@@ 1. Introduction
Monoid expansions provide systematic ways to build a new monoid from a given input monoid, often by inflating the set of idempotents in some structured manner. They were introduced in [1] to facilitate the application of techniques in finite semigroup theory to infinite monoids and semigroups. Most well-studied examples of expansions have the property that when restricting the input to an interesting class of monoids (e.g. groups, cancellative monoids, unipotent monoids), the output of the expansion is a monoid with desirable structural properties. For example, the Szendrei expansion [18] maps: groups to inverse monoids [1], right cancellative monoids to left ample monoids [5] and unipotent monoids to (what are now called) left restriction monoids [6]. In fact, [1] used an earlier formulation, known as the prefix expansion, which is equivalent to the Szendrei expansion in the case the monoid is a group; [14] investigates the prefix expansion of an arbitrary monoid. The underlying construction used in these expansions is a semidirect product formed from the monoid $M$ acting on a particular semilattice of subsets of $M$. It is then natural to consider the ‘parent’ expansion where the semilattice contains all subsets of $M$, and this is our focus here. Specifically, this paper considers the expansion formed via the action of a monoid $M$ on its power set $\mathcal{P}(M)$ viewed as a monoid (indeed, a semilattice) with respect to the operation of union, and with identity element $\emptyset$, with left action of $M$ on $\mathcal{P}(M)$ given by $(m,X)\mapsto mX=\{mx:x\in X\}$ for all $X\in\mathcal{P}(M)$ and all $m\in M$. Throughout the paper we will write $\mathcal{S}(M)$ to denote $\mathcal{P}(M)\rtimes M$ with product given by $(X,x)(Y,y)=(X\cup xY,xy)$, and with identity element $(\emptyset,1)$, where $1$ is the identity element of $M$. We refer to monoids of this form as ‘subset expansions’ to indicate that this construction involves the action of $M$ on its subsets.
@@
Our interest in subset expansions comes from several directions. First, $\mathcal{S}(M)$ contains the Szendrei expansion of $M$ as a monoid subsemigroup, and (as with the Szendrei expansion) if $G$ is a group then $\mathcal{S}(G)$ is an inverse monoid. If $M=F_{X}$ is the free group on a set $X$, we remark that the monoid $\mathcal{S}(F_{X})$ contains: the free inverse, free ample and free left ample monoids on the same set $X$. For a monoid $M$ which is not a group, the semidirect product $\mathcal{S}(M)$ is no longer inverse (indeed, it is not even regular; see Lemma 3.3 below for details). It is then natural to ask under which circumstances $\mathcal{S}(M)$ belongs to one of several more general classes of monoids (e.g. left/right restriction, ample, adequate, abundant).
@@
Second, we note that any inverse monoid satisfies several finitary conditions (that is, conditions satisfied by every finite monoid), including weak left (respectively, right) coherence [2]. A monoid $M$ is said to be weakly left (respectively, right) coherent if every finitely generated left (respectively, right) ideal of $M$ is finitely presented as a left (respectively, right) $M$-act. A classical result of Wheeler [19], interpreted for acts in [9], says that a monoid is left coherent if and only if the class of existentially closed left acts is axiomatisable. Inspired by this, weak left coherency for monoids, and the related finitary conditions we consider here, have also been formulated in terms of the model-theoretic notion of axiomatisability: see [8, 17, 4]. In previous work [13] the construction $\mathcal{S}(M)$ has been utilised to exhibit some interesting examples of monoids that are neither left nor right coherent. We have remarked that the monoid $\mathcal{S}(F_{X})$ contains the free inverse, free ample and free left ample monoids on set $X$, and it is known that each of the latter is weakly left and right coherent [2, 3, 11]; cf. the table in [13]. One is then drawn to ask (see Question 8.6 of [13]): for which monoids $M$ is $\mathcal{S}(M)$ weakly left (respectively right) coherent? It is known that $M$ is weakly left coherent if and only if it is simultaneously left ideal Howson (the intersection of any two finitely generated left ideals is finitely generated) and finitely left equated (the left annihilator congruences of $M$ are finitely generated as left congruences). We initiate the study of finitary conditions in the context of expansions of monoids by examining the behaviour of these two conditions and the closely related conditions of (R) and (L) with respect to the expansion $\mathcal{S}$.
@@
The paper is structured as follows. In Section 2 we outline the preliminary definitions and results required for our paper and in Section 3 we prove some structural results on subset expansions. Our main results are in Section 4. We give precise characterisations of the monoids $M$ for which (R), (L) (Subsection 4.1) or the left and right ideal Howson properties hold in $\mathcal{S}(M)$ (Subsection 4.2). We characterise the monoids $M$ for which $\mathcal{S}(M)$ is finitely left equated. We exhibit a configuration of elements in $M$ that prevent $\mathcal{S}(M)$ from being finitely right equated. We introduce a notion of being strongly finitely right equated and show that $\mathcal{S}(M)$ satisfies this condition if and only if $M$ does. (Subsection 4.3). In particular, we apply our results to demonstrate that if $M$ is a right (left) LCM monoid then $\mathcal{S}(M)$ is weakly right (left) coherent. We provide a number of examples and counterexamples to illustrate and separate the properties under consideration. We finish with some open questions.
@@ 2. Preliminaries
In this section we outline the background necessary for this paper. Where the notions we discuss have clear left-right duals we do not mention them explicitly, except to set up notation. We refer the reader to [15] for a fuller account of the semigroup theoretic notions we discuss. Throughout the paper $M$ denotes a monoid with identity element $1$.
@@ 2.3. A suite of finitary conditions
Another related finitary condition arises as follows. Here the right diagonal act of $M$ is the set $M\times M$, with co-ordinatewise right action by $M$, that is, $(a,b)m=(am,bm)$ for all $(a,b)\in M\times M$ and $m\in M$. A right subact $A$ of $M\times M$ is finitely generated if there is a finite set $X\subseteq A$ such that $A=XM$.
@@
It is easy to see that $\mathbf{R}(a,b)$ is a right subact of the right diagonal act of $M$. We say that $M$ satisfies condition (R) if for all $a,b\in M$ the right subact $\mathbf{R}(a,b)$ is finitely generated.
@@
3. Properties of subset expansions
@@
    $\mathcal{S}(M)$ is always a proper left $\mathcal{E}$-restriction monoid and so certainly left $\mathcal{E}$-Ehresmann; it is left $\mathcal{E}$-ample if and only if $M$ is right cancellative, in which case $\mathcal{E}=E(\mathcal{S}(M))$.
@@
We note that Lemma 3.3 together with Proposition 3.4 yields that the expansion $\mathcal{S}$ maps groups to proper inverse monoids, unipotent monoids to proper left restriction monoids, and right cancellative monoids to left ample monoids. Similarly, by Proposition 3.5 we have that $\mathcal{S}$ maps right abundant monoids to right abundant monoids and left cancellative monoids to right adequate monoids.
@@ 4.2. The ideal Howson properties
The expansion $\mathcal{S}(M)$ of a monoid $M$ is right ideal Howson if and only if $M$ is strongly right ideal Howson. Moreover, $\mathcal{S}(M)$ is principally right ideal Howson if and only if $M$ is principally right ideal Howson.
@@
The expansion $\mathcal{S}(M)$ of a monoid $M$ is left ideal Howson if and only if $M$ is left co-ordinated. Further, $\mathcal{S}(M)$ is principally left ideal Howson if and only if $M$ is left $1$-co-ordinated.
@@
Let $M$ be a right cancellative monoid. The expansion $\mathcal{S}(M)$ is left ideal Howson if and only if $M$ is left ideal Howson.
@@
If $M$ is a left (1-)co-ordinated monoid, and $T$ is a retract of $M$, then $T$ is also left (1-)co-ordinated.
@@ 4.3. The finitely equated conditions
Since $M$ is a retract of $\mathcal{S}(M)$ it follows from [4, Corollary 4.12] that a necessary condition for $\mathcal{S}(M)$ to be finitely right (respectively, left) equated is that $M$ is finitely right (respectively, left) equated, however, as we shall see below, the converse does not hold (in either case). We look first at the case of finitely left equated monoids, providing an exact characterisation of the monoids $M$ for which $\mathcal{S}(M)$ is finitely left equated.
Recalling the left-right dual of Definitions 2.2 and 2.9, we note that for a monoid $M$ and $a\in M$ we have $\mathbf{l}(a)=\{(u,v)\in M\times M:ua=va\}=\mathbf{L}(a,a)$. It therefore makes sense to consider whether $\mathbf{l}(a)$ is finitely generated as a subact of the left diagonal act of $M$, and this is easily seen to imply that $\mathbf{l}(a)$ is finitely generated as a left congruence. (The converse is not true, as we demonstrate in Example 4.18.)
Theorem 4.15.
The expansion $\mathcal{S}(M)$ of a monoid $M$ is finitely left equated if and only if for each $a\in M$ we have that $\mathbf{l}(a)$ is finitely generated as a left $M$-act.
Proof.
Suppose first that for all $a\in M$ we have that $\mathbf{l}(a)$ is finitely generated as a left $M$-act. For an arbitrary $A\subseteq M$ and $a\in M$, we aim to show that $\mathbf{l}((A,a))$ is a finitely generated left congruence. By assumption, $\mathbf{l}(a)$ is finitely generated as an $M$-act; let $X\subseteq\mathbf{l}(a)$ be a finite symmetric generating set, so that $\mathbf{l}(a)=MX$. We claim that $\mathbf{l}((A,a))$ is the left congruence $\lambda_{Y}$ generated by the finite set $Y=\{((\emptyset,1),(A,1))\}\cup\{((pA,q),(qA,p)):(p,q)\in X\}$. To see this, first note that $Y\subseteq\mathbf{l}((A,a))$ since:
$(\emptyset,1)(A,a)$ $=$ $(A,a)=(A,1)(A,a)$
$(pA,q)(A,a)$ $=$ $(pA\cup qA,qa)=(qA\cup pA,pa)=(qA,p)(A,a),$
where the second equality holds for all $(p,q)\in X\subseteq\mathbf{l}(a)$. Now for $((U,u),(V,v))\in\mathbf{l}((A,a))$ we have $U\cup uA=V\cup vA$ and $ua=va$. Since $(u,v)\in\mathbf{l}(a)$, there exist $(p,q)\in X$ and $t\in M$ such that $u=tp$ and $v=tq$. Setting $W=U\cup uA=V\cup vA$ we have:
$(U,u)=(U,u)(\emptyset,1)\,\lambda_{Y}\,(U,u)(A,1)=(W,u)=(W,1)(vA,u)=(W,1)(tqA,tp)=(W,t)(qA,p)$
$(V,v)=(V,v)(\emptyset,1)\,\lambda_{Y}\,(V,v)(A,1)=(W,v)=(W,1)(uA,v)=(W,1)(tpA,tq)=(W,t)(pA,tq),$
and since $(qA,p)\,\lambda_{Y}\,(pA,q)$ this demonstrates that $(U,u)\,\lambda_{Y}\,(V,v)$.
Conversely, suppose that $\mathcal{S}(M)$ is finitely left equated. In particular for all $a\in M$ we have that $\mathbf{l}((\{1\},a)$ is a finitely generated left congruence; let $X$ be a finite symmetric generating set. Now, if $((P,p),(Q,q))\in X$ we have
$(P,p)(\{1\},a)=(P\cup\{p\},pa)=(Q\cup\{q\},qa)=(Q,q)(\{1\},a),$
that is, $P\cup\{p\}=Q\cup\{q\}$ and $(p,q)\in\mathbf{l}(a)$. We claim that $\mathbf{l}(a)=MZ$ where
$Z=\{(1,1)\}\cup\{(p,q):\exists((P,p),(Q,q))\in X\}\subseteq\mathbf{l}(a).$
Since $X$ is finite, $Z$ is necessarily finite. Let $(u,v)\in\mathbf{l}(a).$ If $u=v$ then clearly $(u,v)=u(1,1)$. Suppose then that $u\neq v$. Then
$(\{u\},v)(\{1\},a)=(\{u,v\},va)=(\{u,v\},ua)=(\{v\},u)(\{1\},a),$
demonstrating that $((\{u\},v),(\{v\},u))\in\mathbf{l}(\{1\},a)$. Thus there is an $X$-sequence
$(\{u\},v)=(T_{1},t_{1})(P_{1},p_{1}),(T_{1},t_{1})(Q_{1},q_{1})=(T_{2},t_{2})(P_{2},p_{2}),\ldots,(T_{n},t_{n})(Q_{n},q_{n})=(\{v\},u).$
For $1\leq i\leq n$, set $Z_{i}=T_{i}\cup t_{i}P_{i}$, $z_{i}=t_{i}p_{i}$ and also set $Z_{n+1}=T_{n}\cup t_{n}Q_{n}$ and $z_{n+1}=t_{n}q_{n}=u$. We prove (by induction) that for each $i$ we have $Z_{i}\subseteq\{u,v\}$ and $z_{i}\in\{u,v\}$. In view of our $X$-sequence, this is immediate for $i=1$. Now suppose that $1\leq i\leq n$ is such that $Z_{i}\subseteq\{u,v\}$ and $z_{i}\in\{u,v\}$. Since $((P_{i},p_{i}),(Q_{i},q_{i}))\in X$ we have $P_{i}\cup\{p_{i}\}=Q_{i}\cup\{q_{i}\}$ and so
$\{u,v\}\supseteq Z_{i}\cup\{z_{i}\}=T_{i}\cup t_{i}P_{i}\cup\{t_{i}p_{i}\}=T_{i}\cup t_{i}(P_{i}\cup\{p_{i}\})=T_{i}\cup t_{i}(Q_{i}\cup\{q_{i}\})\supseteq T_{i}\cup t_{i}Q_{i}.$
Thus $Z_{i+1}=T_{i+1}\cup t_{i+1}P_{i+1}=T_{i}\cup t_{i}Q_{i}\subseteq\{u,v\}$ and $z_{i+1}=t_{i+1}p_{i+1}=t_{i}q_{i}\in\{u,v\}$. Now, returning to our $X$-sequence, we have (from the second co-ordinate)
$v=t_{1}p_{1},t_{1}q_{1}=t_{2}p_{2},\ldots,t_{n}q_{n}=u$
where (by the argument above) each term in this sequence, $z_{i}$, is either equal to one of $u$ or $v$. Since $z_{1}=v\neq u=z_{n+1}$, there must exist $i$ such that $z_{i}=t_{i}p_{i}=v$ and $z_{i+1}=t_{i+1}p_{i+1}=t_{i}q_{i}=u$, giving $(u,v)=t_{i}(q_{i},p_{i})\in MZ$ as required. ∎
We state the following immediately consequence of Theorem 4.12 and Theorem 4.15.
Corollary 4.16.
The expansion $\mathcal{S}(M)$ of a monoid $M$ is weakly left coherent if and only if $M$ is left co-ordinated and, for each $a\in M$, $\mathbf{l}(a)$ is finitely generated as a left $M$ act.
Right cancellative monoids provide a particular instance of Theorem 4.15, giving:
Corollary 4.17.
Let $M$ be a right cancellative monoid. Then $\mathcal{S}(M)$ is weakly left coherent if and only if $M$ is left ideal Howson.
Proof.
Suppose $M$ is right cancellative. Then for $a\in M$ we have
$\mathbf{l}(a)=\{(u,v):ua=va\}=\{(u,u):u\in M\}=M(1,1),$
and so by Theorem 4.15, $\mathcal{S}(M)$ is finitely left equated, and so $\mathcal{S}(M)$ is weakly left coherent if and only if $\mathcal{S}(M)$ is left ideal Howson. By Proposition 4.12 this is equivalent to $M$ being left co-ordinated, and since $M$ is right cancellative Proposition 4.9 gives that this is equivalent to the condition that $M$ is left ideal Howson. ∎
As we have alluded to earlier, there exist left abundant monoids $M$ such that $\mathcal{S}(M)$ is not left abundant; making use of Theorem 4.15 we exhibit an example of a semilattice $M$ (which is certainly left abundant) such that $\mathcal{S}(M)$ is not finitely left equated and so certainly not left abundant.
Example 4.18.
Let $M$ be the monoid with presentation as a monoid semilattice given by
$M:=\langle a,b_{i},c_{i},i\in\mathbb{N}\mid b_{i}a=c_{i}a\rangle.$
Since $M$ is a semilattice, it is certainly abundant and so finitely left equated. By construction, we have infinitely many distinct pairs $(b_{i},c_{i})\in\mathbf{l}(a)$, and it is easily seen that $\mathbf{l}(a)$ is not finitely generated as a left $M$-act. By Theorem 4.15, we conclude that $\mathcal{S}(M)$ is not finitely left equated (and in particular not left abundant).
In contrast with the previous example, and turning now to the case of finitely right equated monoids, it can be deduced from Proposition 2.11 and Proposition 3.5 that if $M$ is right abundant, then $\mathcal{S}(M)$ is finitely right equated. We generalise this to a statement about skeletons below.
Definition 4.19.
For a right congruence $\rho_{W}$ on a monoid $M$ generated by a finite set $W$, and for $a,b\in\rho_{W}$, we may call the sequence $c_{1},d_{1},\cdots,c_{n},d_{n}$ appearing in the $W$-sequence in Proposition 2.1 a skeleton from $a$ to $b$ over $W$. We say that $\rho_{W}$ admits a fixed skeleton over $W$ if there is a fixed sequence $c_{1},d_{1},\cdots,c_{n},d_{n}$ that is a skeleton from $a$ to $b$ over $W$ for all $(a,b)\in\rho_{W}$. Further, we say that $M$ is strongly finitely right equated if for all $x\in M$, the right annihilator congruence $\mathbf{r}(x)$ admits a fixed skeleton over some finite generating set.
Notice that several natural classes of monoids are strongly finitely right equated. In particular, every right abundant monoid $M$ is strongly finitely right equated, since for any $x\in M$ and any $a,b\in\mathbf{r}(x)$, taking $e$ to be an idempotent in the $\mathcal{L}^{*}$-class of $x$ we have that $xe=x$ so that $Y:=\{(1,e),(e,1)\}\subseteq\mathbf{r}(a)$ and then for any $(u,v)\in\mathbf{r}(x)$ we see that
u=1u,eu=ev, 1v=vu=1u,\,eu=ev,\,1v=v
is a $Y$-sequence, demonstrating that $\mathbf{r}(x)$ has fixed skeleton $1,e,e,1$. Our next aim is to show that $\mathcal{S}(M)$ is strongly finitely right equated if and only if $M$ is strongly finitely right equated, hence giving a generalisation of the result for right abundance. First notice that the property of being strongly finitely right equated is preserved under taking retracts.
Lemma 4.20.
Let $M$ be a monoid and let $S$ be a retract of $M$. If $M$ is strongly finitely right equated, then so is $S$.
Proof.
Let $\theta:M\rightarrow S$ be a retraction where we consider $S$ as a submonoid of $M$. Fix $a\in S$ and let $\mathbf{r}_{M}(a)$ and $\mathbf{r}_{S}(a)$ denote the right annihilator congruence of $a$ on $M$ and $S$, respectively. By assumption, $\mathbf{r}_{M}(a)$ is strongly finitely right equated and so there is a finite generating set $X$ such that $\mathbf{r}_{M}(a)$ admits the finite skeleton $c_{1},d_{1},\ldots,c_{n},d_{n}$ over $X$. Then, for $(u,v)\in\mathbf{r}_{S}(a)\subseteq\mathbf{r}_{M}(a)$ we have that there exist $t_{1},\ldots,t_{n}\in M$ such that $u=c_{1}t_{1},\,d_{1}t_{1}=c_{2}d_{2},\,\ldots,\,d_{n}t_{n}=v$. Applying the morphism $\theta$ (which fixes each element of $S$) then gives $u=c_{1}\theta t_{1}\theta,\,\,d_{1}\theta t_{1}\theta=c_{2}d_{2}\theta,\,\ldots,\,d_{n}\theta t_{n}\theta=v$. Thus, noting that if $ap=aq$ for $p,q\in M$ we also have $a(p\theta)=a(q\theta)$ where $p,q\in S$, we may take $Y=\{(p\theta,q\theta):(p,q)\in X\}$ and sequence $c_{1}\theta,d_{1}\theta,\ldots,c_{n}\theta,d_{n}\theta$ to see that $S$ is strongly finitely right equated. ∎
Theorem 4.21.
The expansion $\mathcal{S}(M)$ of a monoid $M$ is strongly finitely right equated if and only if $M$ is strongly finitely right equated.
Proof.
One direction follows immediately from Lemma 4.20 since $M$ is a retract of $\mathcal{S}(M)$.